Shift report 20000806 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Aug 06 2000 - 08:18:40 EDT

  • Next message: Konstantin Olchanski: "Re: Shift report 20000805 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Aug  6 08:18:40 2000, Shift: 00:00-08:00
    Supervisor: PS, with: 
    Log book#4 pages 8 - 11
    
    Shift Summary:
    ------------------------------------------------------------
    1. Stable beam from 2:00 to 3:00 at medium intensity. 
       Beam was lost at 3:00 because of "quench interlock 
       in blue ring". At 6:00 exper come in.
    
    2. Run #2063. About 230 "fresh" collisions have been collected.. 
       Mapping for 2 T3 module checked and sent to KO.
    
    3. T2 readout is broken.
       FFS1440 does not respond.
       Konstantin was after midnight and made diagnosis:
       the power on mainframe has to be cycled.
       
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 06 2000 - 08:21:48 EDT